OpenStudy (anonymous):

find the values of x if f(x)=g(x) f(x)=x to the fourth - 2x to the 2nd g(x)= 2x to the 2nd please explain!!

OpenStudy (anonymous):

set \[x^4-2x^2=2x^2\]and solve for x

OpenStudy (anonymous):

start with \[x^4-4x^2=0\] \[x^2(x^2-4)=0\] \[x^2(x+2)(x-2)=0\] \[x=0,x=-2,x=2\]

OpenStudy (anonymous):

so you factor it twice? and its the opposite of the the 2s in the parenthesis?

OpenStudy (anonymous):

yes factor all the way down

OpenStudy (anonymous):

then set each factor = 0 and solve

OpenStudy (anonymous):

ok thanks that helped a lot!